OpenStudy (styxer):

Let xn be the n-th term of the sequence: 1/1 , 1/4 , 3/4 , 1/9 , 3/9 , 5/9 , 1/16 , 3/16 , 5/16 , 7/16 , 1/25... Find the maximal n satisfying xn ≥ 1/10

OpenStudy (templeguy):

I'll answer yours if you answer mine

OpenStudy (templeguy):

thanks

OpenStudy (templeguy):

wait, forget it, you're just gonna say because the brain said so

OpenStudy (itz_sid):

do you mean\[x_n \leftarrow or\rightarrow xn \]

OpenStudy (styxer):

@iTz_Sid I had a little fight with LaTeX and couldn't do anything... It's the first one.

OpenStudy (sooobored):

weird sequence the numerator are increasing prime and denominator are squares

OpenStudy (styxer):

@sooobored yes... The pattern is (on the denominator) 1²,2²,3²,and so on... On the numerator is 1,3,5,7... , and the number of terms that appears is the same of the denominator without being squared. The sum of the numbers with the same denominator will always result in 1.

OpenStudy (caozeyuan):

the problem is that there is no general formula for Xn as a function of n , not that I can think of

OpenStudy (loser66):

I observed it and realized that \(1,\\ (1/4)+(3/4) =1,\\ (1/9)+(3/9)+(5/9)=1,\\ (1/16)+(3/16)+(5/16)+(7/16)=1\cdots\)

OpenStudy (loser66):

So, the second group has the form of \(x_1 + x_2=1\) the third group has the form \(x_1+x_2+x_3=1\) The forth group is \(x_1+x_2+x_3+x_4=1\) We can assume that \(1=\sum_{i=1}^n x_i\)

OpenStudy (loser66):

But I don't get the question!! is it \(x_n\geq\dfrac{1}{10}\) or \(x*n \geq \dfrac{1}{10}\)

OpenStudy (loser66):

The denominators in the group change as a sequence of 1, 4, 9, 16, 25 ,... We can find out the rule of it. It is \(d_n= n^2+2n+1\) (Hopefully you know how to do it)

OpenStudy (loser66):

For the numerator, we can see that on the sequence 1,3,5,7,9,11,... \(S_1= 1\\S_2= 1+3\\S_3=1+3+5\\S_4=1+3+5+7\) so, for "group" of the sequence, the numerator has the form \(N_n= S_n\)

OpenStudy (styxer):

@loser66 the question is a "xn" term... I couldn't do it on latex ;-;

OpenStudy (loser66):

And the sum is = the denominator

OpenStudy (loser66):

We have the formula for it. \(\sum_{i=1}^n odd=n^2\) So far, we have the formula for any group of the sequence is \(\dfrac{n^2}{n^2+2n+1}\)

OpenStudy (loser66):

However, I am too lazy to derive more. hahaha... You can say that I am suddenly stupid.